OpenStudy (anonymous):

can some help me with 3/15 + 2/x

OpenStudy (anonymous):

no because there is a plus not en equal sign

OpenStudy (abb0t):

Find a common denominator. For this example, your common denominator is: 15x That might help guide you.

OpenStudy (anonymous):

what do i do with the common denominator would i get 3+2/15x?

OpenStudy (abb0t):

Well, you know: \[\frac{ 2 }{ a } + \frac{ 1 }{ a } = \frac{ 2 + 1 }{ a } = \frac{ 3 }{ a }\]

OpenStudy (abb0t):

Also, \[\frac{ 1 }{ b } + \frac{ 3 }{ d } = \frac{ 1d + 3b }{ bd }\]

OpenStudy (anonymous):

so my final answer would be 5/15x?

OpenStudy (abb0t):

Not quite. Because in order to get a 15x, you must multiply 3 by x and 2 by 15. \[\frac{ 3 }{ 15 } + \frac{ 2 }{ x } = (\frac{ 3x }{ 15x }) + (\frac{ 2(15) }{ 15x })\]

OpenStudy (abb0t):

The next step is to combine everything into one whole fraction.